$24,000 is deposited for 5 years in an account earning 7% interest. (Round your answers to two decimal places.)

Calculate the interest earned if interest is compounded quarterly.

Answer by Theo(13342)

formula is f = p * (1 + r) ^ n

f is the future value.
p is the present value.
r is the interest rate per time period.
n is the number of time periods.

in your problem:
f is what you want to find.
p is 24,000
r is 7/5 per year divided by 100 = .07 divided by 4 quarters per year = .07/4 = .0175 per quarter.
n is 5 years * 4 quarters per year = 20 quarters.

formula becomes f = 24,000 * (1 + .0175) ^ 20 which becomes f = 33,954.6767

the interest earned is f - p = 33,954.6767 - 24,000 = 9,954.676698.

round to two decimal laces and the solution is 9,954.68.